what is the perimeter of a rectangle with the length of 2 cm and a width of 3 cm
Answer by jim_thompson5910(35256)

Perimeter = 2*(Length + Width)

P = 2*(L+W)

P = 2*(2+3)

P = 2*5

P = 10


So the perimeter is 10 cm